A box of mass 2.00kg is released from rest at the top of a plane inclined 15.3º above horizontal. The coefficient of kinetic friction is 0.22. What will be the speed of the mass after sliding 4.2 m along the plane? Round to the nearest hundredth.



1
Expert's answer
2022-01-31T09:05:18-0500

1) Determine the acceleration:


"a=g(\sin\theta+\mu\cos\theta)."


2) Find the speed:


"v=\sqrt{2ad}=\sqrt{2dg(\sin\theta+\mu\cos\theta)},\\
v=6.26\text{ m/s}."